# Distance-Preserving Spatial Representations in Genomic Data

### v2

@@ 4.4 Theoretical Property
We briefly comment on our finding on the theoretical connection between the proposed distance-preserving loss and the distortion [8] or bi-Lipshitz condition[36, 47]. Specifically, the enforcement result can be viewed as a probably approximately correct (PAC) learning version of the two, summarized in the following theorem.
@@
In Theorem 1, expression (15) resembles the distortion or bi-Lipschitz condition defined in prior works, except that it only holds probabilistically. This probability is incurred by the underlying data generation process and the learned VAE generative model. The distortion constant or bi-Lipschitz constant $L$ is porportional to the expected value of the distance-preserving loss $\mathcal{L}_{\mathrm{DP}}$, which in the ideal case (i.e. perfect generalization) scales with the value of its training error. This means that the encoder network becomes “smoother” as training progresses, and the distance-preserving loss is what regulates such behavior of this generative model. We note that this result summarizes and can be potentially generalized to characterize the theoretical properties of the generative networks regulated with distance-preserving loss, as it is not unique to the dp-VAE framework we are using in this paper. We provide the exact bound to Theorem 1 in the appendix.
@@ 5 Experiments
In this section, we conduct experiments to validate the effectiveness of our method and shed light on some of our method’s practical implications, including training stability and in-sample behavior of dp-VAE (Section 5.2), its comparison against vanilla VAE in out-of-sample evaluation (Section 5.3) and out-of-distribution inference applications, i.e., transfer learning (Section 5.4). We use a total of 27 processed Visium Spatial Gene Expression data from the 10x Genomics database provided by Scanpy [51]. We only select the top $100$ most variable genes for all datasets to ensure model stability and learning feasibility  [^3] Most datasets only contain roughly around three thousand entries, making the problem high-dimensional and numerically unstable. The names of the datasets are the initials of the datasets, whose full names and correspondence can be found in the appendix. Before we proceed, we first describe our experiment setups in the next subsection.
@@ 5.1 Experiment Setups
Model description. Since the VAE model is a special case of dp-VAE when the distance-preserving regularization coefficient is set to zero. Therefore, we will only describe our set-up for the dp-VAE model. The dp-VAE model consists of an encoder and a decoder network, each comprising a one-layer fully connected neural network with $124$ hidden layers and a latent dimension of $2$; therefore, the procedure of solving for an optimization problem can be bypassed by directly using the latent representations as the estimated spatial coordinates. The prior, posterior, and likelihood distributions are all specified as Gaussian distributions. We apply no mask between any pair of entries. The regularization coefficients $\alpha_{1},\beta$ for the ELBO objective are set to one throughout all experiments, therefore for simplicity, we will only refer to $\alpha_{2}$ as the regularization coefficient in our experiment section. The learning rate is set to $5\times 10^{-4}$ or $10^{-4}$ depending on the experiment, and the total number of training epochs is set to $2000$. All experiments are run on NVIDIA GeForce RTX 4070Ti GPU.
Figure 11: The matrix of out-of-distribution spatial reconstruction error incurred by dp-VAE. The model uses each column data as the training data to infer four test datasets labeled on the rows.
Evaluation protocals. We evaluate our model based on the Procrustes distance [15, 16, 25] between the reconstructed and the ground-truth spatial information matrix. Formally, for two datasets denoted by matrices $\textbf{D}_{1},\textbf{D}_{2}\in\mathbb{R}^{n\times m}$, the procedure of calculating the Proscrutes distance between $\textbf{D}_{1}$ and $\textbf{D}_{2}$ can be defined as the procedure of finding the optimal translation, rotation, and scaling transformation from $\mathbb{R}^{m}$ to $\mathbb{R}^{m}$, such that the Frobenius norm between the two matrices is minimized. Formally, it can be described as
@@
Here $\textbf{A}\in\mathbb{R}^{m,m}$ is an orthogonal matrix, $b\in\mathbb{R}^{m}$, and c is a scalar. We will refer to it as the spatial reconstruction error in the following sections when clear from context.
@@ 5.2 In-Sample Training Behaviors
Figure 7 shows the normalized training loss curve of the dp-VAE model with different choices of the regularization coefficient $\alpha_{2}$ averaged over 27 datasets used as training data. Our results show that increasing $\alpha_{2}$ may introduce small abrupt oscillations in the loss curve, globally the loss manages to converge at the same point for all choices of $\alpha_{2}$. This means that different choices of $\alpha_{2}$ do not significantly impact the convergence speed and stability of the training, making its training behavior almost identical to the standard VAE model. In practice, dp-VAE can be tuned and trained very similarly to a standard VAE. However, it is noted that using large learning rates may amplify the oscillation in the training, making it numerically unstable.
@@ 5.3 Out-of-Sample Evaluation
To evaluate the out-of-sample generalization ability and verify that the dp-VAE does not simply memorize the training data, we carry out comparative studies comparing VAE and dp-VAE, as well as different choices of the regularization coefficient $\alpha_{2}$ on out-of-sample spatial reconstruction evaluation tasks. We obtain our training and test data by dividing each one of $27$ datasets by a ratio of $80\%-20\%$, yielding 27 training datasets paired with 27 test datasets.
@@ 5.4 Out-of-Distribution Evaluation and Inference
Figure 11 visualize the normalized spatial reconstruction error matrix, where the columns are the test data where the error is calculated, and the rows are the training data where the dp-VAE model is trained. It can be seen that the train-test pairs VMBSPS2–VMBSP, VMBSPS2–VMBSP, and VMBSA–VMBSAS2 incur relatively small error scales compared to other train-test pairs. This may indicate that the joint distribution of gene expression and spatial context between these sets of tissues are similar, and this shared information may be used to infer one another.
@@ 6 Conclusion and Discussion
In this paper, we proposed a novel representation learning framework based on a variational autoencoder structure to enable tissue spatial reconstruction and imputation from gene expression data. Note that our approach relies minimally on spatial transcriptomics in the sense that they are only used during the training stage as an auxiliary variable, which allows for diverse downstream applications that rely on having spatial context to be carried out when such information is unavailable during the inference stage, opening up the possibility of a broad spectrum of genomics-related studies that might be once unavailable on many datasets. The technical novelty lies in the distance-preserving loss that we introduce to regulate the training process, which helps the model to effectively capture gene expression signals oriented towards describing its spatial context. Spatial reconstruction can be achieved by solving an optimization problem or automatically achieved when the latent space dimension is set to be the dimension of the spatial context. We also point out the theoretical gap between the distance-preserving loss or bi-Lipschitz condition in the context of generative modeling. Finally, we study the training robustness, out-of-sample inference, and transfer learning performance of our proposed method on 27 spatial transcriptomics datasets to evaluate and demonstrate its empirical effectiveness.
A potential future research direction is to more extensively evaluate the method in more challenging transfer learning settings, such as examining the performance of the method in inferring the spatial context of gene expressions that are at a larger scale or with higher spatial resolution, studying the effect of different masking procedures [6], performance in different downstream analysis when integrated with other algorithms, or examine the performance of the method when using reference datasets with higher dissimilarity to the inference datasets. This comparative study might need to incorporate more biological domain expert insights and is more resource-demanding. Nevertheless, we believe that this general framework has strong potential to be modified to various tailored settings.
@@
For any given distribution defined on some space $\mathcal{X}\subseteq\mathbb{R}^{n}$, given arbitrary $10<\delta<$ 1, there exist some $0<M_{1}<M_{2}<+\infty$, such that:
@@
By definition, given any $10<\delta<$ 1, there exists some $R^{\prime}>0$ such that:
@@
$\mathcal{A}$ =$=\left\{M_{1}\leq\|s-s^{\prime}\|\leq M_{2}\right\},$
$\mathcal{B}$ =$=\left\{\lambda\|s-s^{\prime}\|-C\leq\|z-z^{\prime}\|\leq\lambda\|s-s^{\prime}\|+C\right\}$
$\mathcal{C}$ =$=\left\{\lambda\|s-s^{\prime}\|\leq\|z-z^{\prime}\|\leq L\cdot\lambda\|s-s^{\prime}\|\right\},$
we will set $L=M_{2}/M_{1}+C/(\lambda M_{1})$ for now, which then by the definition of the three events, there is
@@
and the fourth line uses (18). Set $C=\mathbb{E}\left[\mathcal{L}_{\textrm{DP}}\right]/(\epsilon-\delta)$ and plug it back to the formula of $L$, then we can organize the previous expression and into concluding when:
@@ Appendix B Additional Related Works
Spatial Data Modeling Autoencoders have been widely utilized in the analysis of spatial data due to their powerful representation learning capabilities. Traditional autoencoders (AE) have been explored in several studies [29, 48]. For instance, SEDR [53] employs a deep auto-encoder network to learn gene representations while incorporating a variational graph auto-encoder to embed spatial information simultaneously, highlighting the integration of spatial and genetic data. The combination of autoencoders with graph neural networks (GNNs) has also proven effective, as shown in works by [31, 60, 17, 30, 59], where such models enhance spatial data representation by leveraging graph structures. Additionally, advanced probabilistic models like the hidden Markov random field have been applied to spatial data, as demonstrated by Giotto [11], which identifies spatial domains using an HMRF model with a spatial neighbor prior. This approach has been extended in studies like [55], further showcasing the potential of probabilistic methods in spatial data analysis.
@@
Spatial Variational Autoencoders Compared to the aforementioned approaches, variational autoencoders (VAEs) are recently attracting more attention in representation learning of single-cell data analysis [12, 13]. They are preferred over traditional autoencoders because they provide a structured, continuous, and interpretable latent space that facilitates robust generative capabilities and handles uncertainty by explicitly modeling distributions. This results in more meaningful data generation and easier manipulation of latent features compared to traditional autoencoders, which can then be used for data integration [12], trajectory inference [13], etc. Furthermore, the study of VAEs lays the groundwork for extending methodologies to other variational inference-based models, such as diffusion models and variational graph autoencoders.
@@
A series of works have been conducted to study the interplay of VAEs with spatial information in the data, and different variants of spatial VAEs have been proposed. [3] enables learning VAE models of images that separate latent variables encoding content from rotation and translation. This can be done by directly constraining them as part of the latent space to be learned. [49] proposed to instead directly modify the latent space from the standard multi-variate Gaussian distributions to low-ranked matrix-variate normal distributions to allow for the better capturing ability of the spatial information present in image data. [1] considered using VAE to learn disentangled representations of spatial networks via careful designs of the model structure and latent variable factorization. They proposed a tractable optimization algorithm to carry out variational learning. More recently, [42] considered the setting of learning spatial Gaussian priors via variational autoencoder. [21] combined graph variational autoencoder with contrastive learning.
@@ Appendix C ELBO Derivation
Below, we present the derivation of lower bound for the log probability density function (PDF) of VAE, namely the evidence lower bound (ELBO). To begin with, the marginal probability of data $x$ can be written as
[…]
where $z$ is a latent random variable. This integral has no closed form and can be usually estimated by Monte Carlo integration with importance sampling, ie,
[…]
Here $q(z|x)$ is the proposed variational distribution, where we can draw sample $z$ from this distribution given $x$ and $\mathbf{a}$. Therefore, by Jensen’s inequality, we can find evidence lower bound (ELBO) of the conditional PDF:
@@ Appendix D KL Divergence
$=\frac{1}{2}\log\frac{|\bm{\Sigma}_{2}|}{|\bm{\Sigma}_{1}|}-\frac{1}{2}\mathop{\mathrm{tr}}\ \{I_{d}\}+\frac{1}{2}(\bm{\mu}_{1}-\bm{\mu}_{2})^{T}\bm{\Sigma}_{2}^{-1}(\bm{\mu}_{1}-\bm{\mu}_{2})$
$+\frac{1}{2}\mathop{\mathrm{tr}}\{\bm{\Sigma}_{2}^{-1}\bm{\Sigma}_{1}\}$
@@
Appendix E Additional Experiment Results
In this section, we show an additional earlier version of experiment result evaluating the spatial autocorrelation and isometry preserving property of our proposed dp-VAE model. Below we introduce the two main evaluation metrics used.
  • Moran’s I [28] is a correlation coefficient that measures the overall spatial autocorrelation of a dataset. Intuitively, for a given gene, it measures how one spot is similar to others surrounding it. If the spots are attracted (or repelled) by each other, it implies they are not independent. Thus, the presence of autocorrelation indicates the spatial pattern of gene expression. Moran’s I value ranges from –1 to 1, where a value close to 1 indicates a clear spatial pattern, a value close to 0 indicates random spatial expression, and a value close to –1 indicates a chessboard-like pattern. To evaluate the spatial variability of a given gene, we calculate the Moran’s I using the following formula,

### v3

@@ 4.4 Theoretical Property
We briefly comment on our findings on the theoretical connection between the proposed distance-preserving loss and the distortion [chennuru2018measures] or bi-Lipshitz condition[mahabadi2018nonlinear, verine2023expressivity]. Specifically, the enforcement result can be viewed as a probably approximately correct (PAC) learning version of the two, summarized in the following theorem.
@@
In Theorem 1, expression (9) resembles the distortion or bi-Lipschitz condition defined in prior works, except that it only holds probabilistically. This probability is incurred by the underlying data generation process and the learned VAE generative model. The distortion constant or bi-Lipschitz constant $L$ is proportional to the expected value of the distance-preserving loss $\mathcal{L}_{\mathrm{DP}}$, which in the ideal case (i.e. perfect generalization) scales with the value of its training error. This means that the encoder network becomes “smoother” as training progresses, and the distance-preserving loss is what regulates such behavior of this generative model. We note that this result summarizes and can be potentially generalized to characterize the theoretical properties of the generative networks regulated with distance-preserving loss, as it is not unique to the dp-VAE framework we are using in this paper. We provide the exact bound to Theorem 1 in the appendix.
@@ 5 Experiments
In this section, we conduct experiments to validate the effectiveness of our method and shed light on some of our method’s practical implications, including training stability and in-sample behavior of dp-VAE (Section 5.2), its comparison against vanilla VAE in out-of-sample evaluation (Section 5.3), and out-of-distribution inference applications, i.e., transfer learning (Section 5.4). We use a total of 27 processed Visium Spatial Gene Expression data from the 10x Genomics database provided by Scanpy [wolf2018scanpy]. We only select the top $100$ most variable genes for all datasets to ensure model stability and learning feasibility  [^3] Most datasets only contain roughly around three thousand entries, making the problem high-dimensional and numerically unstable. The names of the datasets are the initials of the datasets, whose full names and correspondence can be found in the appendix. Before we proceed, we first describe our experiment setups in the next subsection.
@@ 5.1 Experiment Setups
Model description. Since the VAE model is a special case of dp-VAE when the distance-preserving regularization coefficient is set to zero. Therefore, we will only describe our setup for the dp-VAE model. The dp-VAE model consists of an encoder and a decoder network, each comprising a one-layer fully connected neural network with $128$ hidden layers and a latent dimension of $2$; therefore, the procedure of solving for an optimization problem can be bypassed by directly using the latent representations as the estimated spatial coordinates. The prior, posterior, and likelihood distributions are all specified as Gaussian distributions. We apply no mask between any pair of entries. The regularization coefficients $\alpha_{1},\beta$ for the ELBO objective are set to one throughout all experiments; therefore, for simplicity, we will only refer to $\alpha_{2}$ as the regularization coefficient in our experiment section. The learning rate is set to $5\times 10^{-4}$ or $10^{-4}$ depending on the experiment, and the total number of training epochs is set to $2000$. All experiments are run on an NVIDIA GeForce RTX 4070Ti GPU.
Figure 10: The matrix of out-of-distribution spatial reconstruction error incurred by dp-VAE. The model uses each column of data as the training data to infer four test datasets labeled on the rows.
Evaluation protocals. We evaluate our model based on the Procrustes distance [goodall1991procrustes, gower1975generalized, krzanowski2000principles] between the reconstructed and the ground-truth spatial information matrix. Formally, for two datasets denoted by matrices $\textbf{D}_{1},\textbf{D}_{2}\in\mathbb{R}^{n\times m}$, the procedure of calculating the Proscrutes distance between $\textbf{D}_{1}$ and $\textbf{D}_{2}$ can be defined as the procedure of finding the optimal translation, rotation, and scaling transformation from $\mathbb{R}^{m}$ to $\mathbb{R}^{m}$, such that the Frobenius norm between the two matrices is minimized. Formally, it can be described as
@@
Here $\textbf{A}\in\mathbb{R}^{m,m}$ is an orthogonal matrix, $b\in\mathbb{R}^{m}$, and $c$ is a scalar. We will refer to it as the spatial reconstruction error in the following sections when clear from context. Across all of our experiments, the errors are normalized within $[0,1]$ for brevity of comparison.
@@ 5.2 In-Sample Training Behaviors
Figure 5 shows the normalized training loss curve of the dp-VAE model with different choices of the regularization coefficient $\alpha_{2}$ averaged over 27 datasets used as training data. Our results show that increasing $\alpha_{2}$ may introduce small abrupt oscillations in the loss curve; globally, the loss manages to converge at the same point for all choices of $\alpha_{2}$. This means that different choices of $\alpha_{2}$ do not significantly impact the convergence speed and stability of the training, making its training behavior almost identical to the standard VAE model. In practice, dp-VAE can be tuned and trained very similarly to a standard VAE. However, it is noted that using large learning rates may amplify the oscillation in the training, making it numerically unstable.
@@ 5.3 Out-of-Sample Evaluation
To evaluate the out-of-sample generalization ability and verify that the dp-VAE does not simply memorize the training data, we carry out comparative studies comparing VAE and dp-VAE, as well as different choices of the regularization coefficient $\alpha_{2}$ on out-of-sample spatial reconstruction evaluation tasks. We obtain our training and test data by dividing each one of $27$ datasets by a ratio of $80\%-20\%$, yielding 27 training datasets paired with 27 test datasets. We note that the comparison between dp-VAE and VAE primarily serves as an ablation study to assess the effectiveness of the proposed distance-preserving loss. This comparison is also fair, as, to the best of our knowledge, no existing model is designed to exploit spatial information $S$ without requiring it at the inference stage, which is a setting unique to spatial reconstruction tasks defined in this paper (Section 3).
@@ 5.4 Out-of-Distribution Evaluation and Inference
Figure 10 visualizes the normalized spatial reconstruction error matrix, where the columns are the test data where the error is calculated, and the rows are the training data where the dp-VAE model is trained. It can be seen that the train-test pairs VMBSPS2–VMBSP, VMBSPS2–VMBSP, and VMBSA–VMBSAS2 incur relatively small error scales compared to other train-test pairs. This may indicate that the joint distribution of gene expression and spatial context between these sets of tissues is similar, and this shared information may be used to infer one another.
@@
Finally, we do want to note that in Figure 10, the off-diagonal elements of the matrix (out-of-distribution error) take values close to one, while the diagonal elements (in-sample error) remain relatively low and close to zero. This result shows that the out-of-distribution inference of our method has a large error relative to its in-sample inference performance. We hypothesize that the large errors are primarily attributable to our zero-shot inference setting, where test data originates directly from different brain sectors than the training data without direct overlap. This setting is particularly challenging because: ($i$) different brain regions exhibit distinct gene expression patterns and spatial organizations, and ($ii$) our model has no prior exposure to the target domain’s characteristics during training. Such performance degradation is expected in cross-domain spatial genomics, as the learned spatial-expression mappings may not generalize across anatomically distinct regions.
Given the inherent challenges of zero-shot learning, one potential way to mitigate this limitation is through fine-tuning. In practical genomic studies, researchers often have access to a small amount of target-domain data (or data closely resembling it), which can be leveraged for fine-tuning. This strategy has been extensively tested in prior work and has consistently shown substantial error reductions. Therefore, it is reasonable to hypothesize that, with suitable fine-tuning, the out-of-distribution inference task in our setting could also be improved. In general, achieving strong zero-shot generalization remains difficult and constitutes an important direction for future research.
@@ 6 Conclusion and Discussion
In this paper, we propose a novel representation learning framework based on a variational autoencoder structure to enable tissue spatial reconstruction and imputation from gene expression data. Note that our approach relies minimally on spatial transcriptomics in the sense that they are only used during the training stage as an auxiliary variable, which allows for diverse downstream applications that rely on having spatial context to be carried out when such information is unavailable during the inference stage, opening up the possibility of a broad spectrum of genomics-related studies that might be once unavailable on many datasets. The technical novelty lies in the distance-preserving loss that we introduce to regulate the training process, which helps the model to effectively capture gene expression signals oriented towards describing its spatial context. Spatial reconstruction can be achieved by solving an optimization problem or automatically achieved when the latent space dimension is set to be the dimension of the spatial context. We also point out the theoretical gap between the distance-preserving loss or bi-Lipschitz condition in the context of generative modeling. Finally, we study the training robustness, out-of-sample inference, and transfer learning performance of our proposed method on 27 spatial transcriptomics datasets to evaluate and demonstrate its empirical effectiveness.
A potential future research direction is to more extensively evaluate the method in more challenging transfer learning settings, such as examining the performance of the method in inferring the spatial context of gene expressions that are at a larger scale or with higher spatial resolution, studying the effect of different masking procedures [chen2022local], performance in different downstream analysis when integrated with other algorithms, or examine the performance of the method when using reference datasets with higher dissimilarity to the inference datasets. This comparative study might need to incorporate more biological domain expert insights and is more resource-demanding. Nevertheless, we believe that this general framework has strong potential to be modified to various tailored settings.
@@
For any given distribution defined on some space $\mathcal{X}\subseteq\mathbb{R}^{n}$, given arbitrary $0<\delta<1$, there exist some $0<M_{1}<M_{2}<+\infty$, such that:
@@
By definition, given any $0<\delta<1$, there exists some $R^{\prime}>0$ such that:
@@
$\mathcal{A}$ $=\left\{M_{1}\leq\|s-s^{\prime}\|\leq M_{2}\right\},$
$\mathcal{B}$ $=\left\{\lambda\|s-s^{\prime}\|-C\leq\|z-z^{\prime}\|\leq\lambda\|s-s^{\prime}\|+C\right\}$
$\mathcal{C}$ $=\left\{\lambda\|s-s^{\prime}\|\leq\|z-z^{\prime}\|\leq L\cdot\lambda\|s-s^{\prime}\|\right\},$
we will set $L=M_{2}/M_{1}+C/(\lambda M_{1})$ for now, which then, by the definition of the three events, there is
@@
and the fourth line uses (18). Set $C=\mathbb{E}\left[\mathcal{L}_{\textrm{DP}}\right]/(\epsilon-\delta)$ and plug it back into the formula of $L$, then we can organize the previous expression and conclude when:
@@ Appendix B Additional Related Works
Spatial Data Modeling Autoencoders have been widely utilized in the analysis of spatial data due to their powerful representation learning capabilities. Traditional autoencoders (AE) have been explored in several studies [li2024high, wan2023integrating]. For instance, SEDR [xu2024unsupervised] employs a deep auto-encoder network to learn gene representations while incorporating a variational graph auto-encoder to embed spatial information simultaneously, highlighting the integration of spatial and genetic data. The combination of autoencoders with graph neural networks (GNNs) has also proven effective, as shown in works by [long2023spatially, zhou2023integrating, guo2023spiral, liu2023comprehensive, zhang2022graph], where such models enhance spatial data representation by leveraging graph structures. Additionally, advanced probabilistic models like the hidden Markov random field have been applied to spatial data, as demonstrated by Giotto [dries2021giotto], which identifies spatial domains using an HMRF model with a spatial neighbor prior. This approach has been extended in studies like [yang2022sc], further showcasing the potential of probabilistic methods in spatial data analysis.
@@
Spatial Variational Autoencoders Compared to the aforementioned approaches, variational autoencoders (VAEs) are recently attracting more attention in representation learning of single-cell data analysis [du2022robust, du2020model]. They are preferred over traditional autoencoders because they provide a structured, continuous, and interpretable latent space that facilitates robust generative capabilities and handles uncertainty by explicitly modeling distributions. This results in more meaningful data generation and easier manipulation of latent features compared to traditional autoencoders, which can then be used for data integration [du2022robust], trajectory inference [du2020model], etc. Furthermore, the study of VAEs lays the groundwork for extending methodologies to other variational inference-based models, such as diffusion models and variational graph autoencoders.
@@
A series of works have been conducted to study the interplay of VAEs with spatial information in the data, and different variants of spatial VAEs have been proposed. [bepler2019explicitly] enables learning VAE models of images that separate latent variables encoding content from rotation and translation. This can be done by directly constraining them as part of the latent space to be learned. [wang2019spatial] proposed to instead directly modify the latent space from the standard multivariate Gaussian distributions to low-ranked matrix-variate normal distributions to allow for better capturing ability of the spatial information present in image data. [guo2021deep] considered using VAE to learn disentangled representations of spatial networks via careful designs of the model structure and latent variable factorization. They proposed a tractable optimization algorithm to carry out variational learning. More recently, [semenova2022priorvae] considered the setting of learning spatial Gaussian priors via a variational autoencoder. [hu2024spatially] combined graph variational autoencoder with contrastive learning.
@@ Appendix C ELBO Derivation
Below, we present the derivation of a lower bound for the log probability density function (PDF) of VAE, namely the evidence lower bound (ELBO). To begin with, the marginal probability of data $x$ can be written as
$\log f_{\theta}(x)=\log\int p_{\theta}(x,z)dz,$
where $z$ is a latent random variable. This integral has no closed form and can usually be estimated by Monte Carlo integration with importance sampling, ie,
$\int p_{\theta}(x,z)dz=\mathbb{E}_{z\sim q(\cdot|x)}\left[\frac{p_{\theta}(x,z)}{q(z|x)}\right].$
Here $q(z|x)$ is the proposed variational distribution, where we can draw a sample $z$ from this distribution given $x$ and $\mathbf{a}$. Therefore, by Jensen’s inequality, we can find the evidence lower bound (ELBO) of the conditional PDF:
@@ Appendix D KL Divergence
$\hskip-21.52771pt=\frac{1}{2}\log\frac{|\bm{\Sigma}_{2}|}{|\bm{\Sigma}_{1}|}-\frac{1}{2}\mathop{\mathrm{tr}}\ \{I_{d}\}+\frac{1}{2}(\bm{\mu}_{1}-\bm{\mu}_{2})^{T}\bm{\Sigma}_{2}^{-1}(\bm{\mu}_{1}-\bm{\mu}_{2})$
$\hskip-8.61108pt+\frac{1}{2}\mathop{\mathrm{tr}}\{\bm{\Sigma}_{2}^{-1}\bm{\Sigma}_{1}\}$
@@
Appendix E Additional Experiment Details
In this section, we show an additional earlier version of the experiment result evaluating the spatial autocorrelation and isometry preserving property of our proposed dp-VAE model. Below, we introduce the two main evaluation metrics used.
  • Moran’s I [li2007beyond] is a correlation coefficient that measures the overall spatial autocorrelation of a dataset. Intuitively, for a given gene, it measures how one spot is similar to others surrounding it. If the spots are attracted (or repelled) by each other, it implies they are not independent. Thus, the presence of autocorrelation indicates the spatial pattern of gene expression. Moran’s I value ranges from –1 to 1, where a value close to 1 indicates a clear spatial pattern, a value close to 0 indicates random spatial expression, and a value close to –1 indicates a chessboard-like pattern. To evaluate the spatial variability of a given gene, we calculate Moran’s I using the following formula,